### v1

# Reviewing Scientific Papers for Critical Problems With Reasoning LLMs: Baseline Approaches and Automatic Evaluation

Abstract
Recent advancements in large language models have sparked interest in utilizing them to assist the peer review process of scientific publication. Instead of having AI models generate reviews in the same way as human reviewers, we propose adopting them as manuscript quality checkers. We introduce several baseline approaches and an extendable automatic evaluation framework using top LLMs as judges to tackle the difficulty of recruiting domain experts for manual evaluation. Utilizing papers withdrawn from arXiv, we validated our proposed methods with several leading reasoning LLMs from different providers and assessed their performance and API costs for identifying critical errors and unsoundness problems. The OpenAI o3 model performed the best, while o4-mini was the most cost-effective one in our evaluation. This paper provides insights into document-based scientific understanding/reasoning and lays the foundation for future applications.
Reviewing Scientific Papers for Critical Problems With Reasoning LLMs: Baseline Approaches and Automatic Evaluation
Tianmai M. Zhang  Neil F. Abernethy University of Washington {tianmai, neila}@uw.edu
1 Introduction
Recent advancements in the domain intelligence of large language models (LLMs) have initiated interest in utilizing them to assist the peer review process of scientific publication, in hopes of reducing the reviewer crisis due to the skyrocketing number of paper submissions in recent years. Researchers have reported receiving reviews that have likely been written with LLMs, and concern is growing over whether and how AI can be responsibly applied to aid peer review (Naddaf, 2025; Kwon, 2025). The extensive knowledge and high efficiency of LLMs seem promising for streamlining laborious peer review, but irresponsible uses of LLM-generated reviews could significantly undermine trust in the long-established peer review process and even publisher credibility.
Several studies have formally explored and evaluated the quality of LLM-generated reviews. An early study by Liang et al. (2024) reported substantial overlap between GPT-generated reviews and human reviews, and participant-reported helpfulness of GPT-generated feedback. Later studies (Du et al., 2024; Zhou et al., 2024; Shin et al., 2025) revealed defects in LLM-generated reviews, such as superficial comments and a lack of criticism or novelty assessment. Other studies (D’Arcy et al., 2024; Gao et al., 2024; Taechoyotin et al., 2024; Tan et al., 2024; Tyser et al., 2024; Yu et al., 2024; Zhu et al., 2025) developed technical methods to improve LLM review generation. However, all of these studies focused on the scenario where LLMs generate full reviews as human reviewers do, and the most common evaluation method was comparing LLM-generated reviews and human reviews, either manually or computationally.
Inspired by the Black Spatula Project [^1] https://the-black-spatula-project.github.io, we propose using LLMs as manuscript quality checkers rather than efficient reviewers. In this way, LLMs are no longer competing with the role of human reviewers but instead completing necessary sub-tasks for them, thereby saving reviewers’ time and allowing them to focus more on leveraging their domain expertise to evaluate important aspects of the manuscript, such as completeness, coherence, novelty, and significance.
In this work, we consider the identification of critical errors and unsoundness problems as the main goal of an LLM-driven manuscript quality checker and present an extensible framework, including several baseline approaches and an automatic evaluation pipeline, to support related tasks. Utilizing papers withdrawn from arXiv, we then validate our proposed methods with several top-performing reasoning LLMs and assess their performance and costs to inform future research and applications.
2 Methods
2.1 Dataset
We utilized WithdrarXiv (Rao et al., 2024), a large-scale dataset of papers withdrawn from arXiv by September 2024, along with associated retraction comments from authors and well-defined retraction categories. The most common retraction category, "factual/methodological/other critical errors in manuscript", contains 6,018 candidate cases with critical errors that would potentially invalidate study conclusions, such as flawed experimental designs, incorrect data analyses, and proof/lemma errors (Rao et al., 2024).
Since retraction comments were typically short and did not necessarily contain clear mentions of errors (e.g., "This paper has been withdrawn due to some mistakes"), we further filtered the dataset with the help of LLMs. Specifically, de-identified retraction comments were first provided to Gemini 2.5 Flash (preview-04-17) to determine whether each retraction comment clearly specified the error. This step resulted in a subset of 2,190 cases. Our manual review further excluded cases that (1) were incorrectly identified during LLM screening; (2) belonged to different versions of the same paper; (3) were not in English; (4) suspiciously used exactly the same retraction reason ("The author has withdrawn this paper due to a crucial sign error in equation 1") that seems to be a template in early years; (5) contained problems that were unlikely to be detectable from the manuscript alone (e.g., an error in a core reference, a bug in code, unreliable data, problems revealed by new observations, etc.). We also corrected mistakenly redacted theorem names in the retraction comments. The final dataset, named WithdrarXiv-Check, contains 1,225 cases.
We randomly sampled 20% of the dataset (245 cases) as the test set for evaluation experiments. The remaining 80% (980 cases) of the dataset was set aside for training and validation, although these steps were not considered in this work whose main objective was to establish baseline approaches and evaluation methods. Dataset characteristics are provided in Table 1.
Train Test
Sample size 980 245
Time span, n (%)
   2007–2012 155 (16%) 32 (13%)
   2013–2018 487 (50%) 114 (47%)
   2019–2024 338 (34%) 99 (40%)
Main subject, n (%)
   Math 492 (50%) 128 (52%)
   Physics [^2] Contains physics, cond-mat, quant-ph, astro-ph, etc. 256 (26%) 70 (29%)
   Computer Science 196 (20%) 37 (15%)
   Others [^3] Contains stat, q-bio, eess, econ, and q-fin. 36 (4%) 10 (4%)
Page count
   Median 14 14
   [Min, Max] [1, 156] [2, 136]
LaTeX script 216 (88%)
available, n (%)
Table 1: Dataset characteristics.
2.2 Baseline Approaches
We propose three baselines for LLMs to perform quality checks based on different approaches to ingest the papers: (1) paper PDF as an attachment, (2) optical character recognition (OCR) results of the PDF in the prompt, and (3) LaTeX script of the paper in the prompt. The first approach simulates the common real-world scenario wherein users upload the PDF file of a paper as an attachment and ask an LLM to read it. However, the PDF preprocessing pipeline (and its performance) may vary by LLM vendor.
The other two approaches based on OCR or LaTeX help decouple PDF preprocessing steps from LLM inference by providing the same input to all models, thereby enabling fairer comparisons of their competencies. However, LLM performance under the OCR approach could be limited considerably by OCR errors, especially when these errors appear in key math formulae, which are potentially more difficult to accurately transcribe. The LaTeX-based approach could perfectly retain the original math formulae, albeit with the inclusion of LaTeX markup, but may also introduce additional noise and information loss (e.g., page and section number). This method is also restricted to papers with available LaTeX source scripts.
Considering the nature of the dataset, we evaluated the PDF-based approach and the LaTeX-based approach in this work. For the small proportion of papers without available LaTeX scripts (Table 1), we resorted to the problems identified by the same model through the PDF-based approach. Images were ignored when using LaTeX scripts in the absence of a simple method to provide LLMs with images and correct image referencing at the same time. For future work on papers in other scientific domains where PDF is the standard format for paper dissemination, we recommend testing the OCR-based approach as one of the baselines.
In our experiments, both approaches utilized the same simplistic, general task instruction (Appendix A). In short, LLMs were instructed to produce a list of up to $k$ problems or errors that are the most critical in a given paper. The prompt was not customized for our dataset that is rich in math and physics papers. Each LLM quality checker was tested $n_{c}$ ($c$ for checker) times with each paper in consideration of potential variations in outputs.
2.3 Evaluation
Considering the daunting cost of recruiting domain experts to manually evaluate LLM-identified scientific errors, we propose an automatic evaluation pipeline to streamline the process. Inspired by LLM-as-a-judge and the idea of LLMs - You Can’t Please Them All [^4] https://www.kaggle.com/competitions/llms-you-cant-please-them-all, we utilized $m$ top-performing LLMs to serve as judges. Ideally, LLM judges should be from different LLM vendors to maximize preference diversity. Each LLM judge independently evaluates an LLM quality checker’s problem submissions one by one for $n_{j}$ ($j$ for judge) times to determine whether they contain an exact match to the gold error description from the authors. For $n_{j}>1$, the most self-consistent answer of the judge is taken as its final decision. If an LLM quality checker receives a majority of (or all, for a stricter evaluation) affirmative votes from LLM judges, it is deemed to have made a hit on the paper. LLM quality checkers were primarily evaluated by their hit rates on test papers. Since hit rate is likely associated with $k$, the number of problems/errors allowed in generation, we report this metric as the Hit Rate at $k$ (HR@$k$). For $n_{c}>1$, i.e., each LLM quality checker is tested on a paper more than once, the metric becomes the Mean Hit Rate at $k$ (MHR@$k$).
2.4 Experiment Setup
In this work, we took $k=5$, $n_{c}=n_{j}=1$, and $m=2$, i.e., each LLM quality checker was tested once with each paper and was allowed to report up to 5 problems, and 2 LLMs served as judges, each judging a problem submission once.
The following reasoning LLMs were tested as paper quality checkers: Google’s Gemini 2.5 Pro (preview-05-06) and Gemini 2.5 Flash (preview-04-17); OpenAI’s o3 (2025-04-16) and o4-mini (2025-04-16); Anthropic’s Claude 3.7 Sonnet (20250219).
The two LLM judges are Gemini 2.5 Pro (preview-05-06) and o3 (2025-04-16). We initially planned for $m=3$ with Claude 3.7 Sonnet as the last judge, but its overly low hit rates under the PDF-based approach (Table 3) indicate that it is unqualified to serve as a judge in this task. Under $m=2$, both judges must vote affirmatively to confirm a hit.
LLMs were accessed via their official APIs using Python. Model parameters used are provided in Appendix B. We set model temperature to 0 and fixed the decoding seed if possible to minimize answer variability, although reasoning models may still have inherent uncertainty. Reasoning effort or thinking budget was kept as the default or automatic setting if applicable. LLMs were not given access to any tools, including web search.
3 Results
Checker Model # Prob. Identified Hit Rate Avg. Token Usage (/paper) Cost Est.
Avg. Q1 Q3 @5 (%) Input Think Output (\$/paper)
PDF as an attachment
Gemini 2.5 Pro 3.3 3 4 55.1 4,678 14,228 881 0.157
Gemini 2.5 Flash 3.8 3 5 53.5 4,678 8,713 644 0.033
o3 (medium) 4.8 5 5 64.9 16,594 3,152 729 0.321
o4-mini (medium) 4.7 5 5 59.6 17,760 3,582 701 0.038
Claude 3.7 Sonnet 1.6 0 4 16.3 43,357 1,630 311 0.159
LaTeX script in prompt
Gemini 2.5 Pro 3.3 3 4 54.7 20,644 18,069 1,033 0.217
Gemini 2.5 Flash 3.6 3 5 50.6 20,644 13,247 667 0.052
o3 (medium) 4.8 5 5 71.0 21,990 3,156 927 0.383
o4-mini (medium) 4.5 5 5 61.6 22,287 3,421 685 0.043
Claude 3.7 Sonnet 3.4 1 5 33.1 28,284 2,701 515 0.133
Table 2: Numbers of identified problems, hit rates, token usage, and estimated costs of LLM quality checkers.
Table 2 provides the main results assessing LLM quality checkers. Besides the number of identified problems and hit rate, we recorded token usage to inform future work that seeks to apply LLMs to check a larger number of papers. Average costs of reviewing a paper under each pipeline-LLM combination were estimated based on API pricing at the time of our experiments in early May 2025.
Compared to the Gemini family, OpenAI’s o-series models tended to make full use of the 5 problem slots allowed and achieved higher estimated hit rates. Considering the high cost of o3, o4-mini seems to be the most cost-effective choice for this task. After switching to LaTeX, hit rates of Gemini models slightly decreased, likely due to a loss of information, e.g. from figures and section numbers. In contrast, the hit rates of o-series models increased. This could suggest that they received specialized training with LaTeX or a larger LaTeX training set, potentially allowing this syntax to invoke their scientific capabilities more effectively. When papers are provided as PDFs, Claude 3.7 Sonnet found no problem in 64.9% of test papers, leading to a low hit rate of 16.3%. Interestingly, both its number of identified problems and hit rate increased after switching to LaTeX, suggesting potential obstacles in Anthropic’s PDF ingestion pipeline.
Large differences in the input token usage between each model family reflect distinct PDF ingestion pipelines of the three LLM vendors. Unlike Gemini models that understand PDFs purely with vision [^5] https://ai.google.dev/gemini-api/docs/document-processing, OpenAI and Anthropic provide their LLMs with both the extracted text and an image of each page [^6] https://platform.openai.com/docs/guides/pdf-files?api-mode=responses, [^7] https://docs.anthropic.com/en/docs/build-with-claude/pdf-support. It remains unclear why Claude has an exceptionally high input token usage for PDF files, and why o4-mini sometimes used slightly more input tokens than o3. After switching to unified LaTeX scripts, all models used the same magnitude of input tokens, although Claude still consumed more. Regarding thinking token usage, Gemini models notably spent several times as many thinking tokens as o3 and o4-mini under both approaches, but this potentially overthinking behavior did not result in higher hit rates.
Additional hit rate results by individual judge are shown in Table 3 in the Appendix. Compared to performance scores determined by a single judge, final scores reasonably dropped due to the difficulty of receiving affirmative votes from both judges, which demonstrates the resistance of our multi-judge approach to inflation in automatic evaluation.
4 Discussion
This work introduces and validates a framework for automatic evaluation of LLMs for scientific quality checking. Evaluation results revealed impressive performance of leading LLMs in reviewing scientific papers for critical errors and unsoundness problems. We believe that our proposed design, after careful improvement and domain expansion, has the potential to serve as a new benchmark for document-based scientific understanding and reasoning. Our task design is also generalizable to the detection of many other types of errors or problems that may appear in scientific papers, such as data errors, content inconsistencies, unmet publication requirements, nomenclature problems, validity restrictions, and undeclared limitations.
The idea of using natural language processing (NLP) for detecting flaws in scientific papers is, in fact, not completely new (Kuznetsov et al., 2024). Previous studies have explored the application of NLP in the detection of statistical reporting inconsistencies (Nuijten and Polanin, 2020), mathematical or conceptual errors (Liu and Shah, 2023), and inaccurate citations (Sarol et al., 2024; Zhang and Abernethy, 2024). There are also attempts in the industry to develop LLM-driven tools for similar purposes (Naddaf, 2025). The core novelty of our work is that we present a reusable framework for the formal evaluation of LLM capabilities in a wide range of tasks related to scientific quality checking using real full papers.
In this work, we only assessed the performance of the most simplistic approach—providing an LLM with a paper and tasking it with finding problems. More complex approaches are likely to demonstrate further gains in LLM performance. For example, one can customize the prompt based on the scientific field the paper belongs to, design a propose-then-verify workflow, fine-tune models for reviewing tasks, or adopt multi-agent collaboration.
One potential concern over the validity of our evaluation results is that LLMs might have seen the test papers (or other versions of them) in their training data. However, recent studies on data contamination and membership inference attacks (Duan et al., 2024; Fu et al., 2025) suggest that LLMs are unlikely to memorize individual instances from pre-training data due to the combination of few training epochs and enormous corpora, let alone the fact that reasoning LLMs experienced complex fine-tuning and post-training processes. Nevertheless, it would be beneficial to evaluate LLMs on papers that are unlikely to be included in training, such as those published after the knowledge cutoff dates of LLMs.
Copyright restrictions pose a key challenge to the automatic review of scientific papers in multiple fields at scale. In this work, we circumvent this challenge by utilizing publicly available arXiv papers. In real-world applications of LLMs, researchers must consider confidentiality and legal risks before sending private, unpublished manuscripts or copyrighted papers to public interfaces provided by LLM service vendors. We also advise developers who wish to apply LLM review at scale to refrain from posting LLM-generated reviews online before performing careful examination for false positives.
Furthermore, we would like to restate our standpoint that human experts should always be at the center of peer review. Although our results demonstrate seemingly impressive capabilities of LLMs in finding critical problems in papers, it should not be interpreted as meaning that LLMs are competent to replace human reviewers. Instead, journal publishers and conference organizers may consider incorporating LLM quality checkers into initial assessments of manuscripts (Bauchner and Rivara, 2024), thereby reducing the reviewer burden. However, any LLM-aided review processes must be implemented ethically and transparently, in cooperation with the wider scientific community to ensure that it results in a net increase in trust in the scientific literature. This is especially important at a time when AI tools are displacing human employees and trust in science is being questioned.
5 Limitation
As a proof-of-concept study, this work has several limitations. First, only closed-source reasoning LLMs were evaluated. Future work can consider comparing different PDF preprocessing pipelines and open-source LLMs. Second, our evaluation metrics may suffer inaccuracies and inflation due to (1) ambiguities in paper authors’ retraction comments (e.g., "withdrawn due to a crucial error in Lemma 2", but what exactly is the error?), (2) automatic evaluation utilizing LLMs without human involvement, and (3) our settings that each LLM quality checker was tested only once per paper and each LLM judge graded each submission only once. Nevertheless, a higher score under parallel evaluation still reasonably indicates better performance. We expect gradual improvements in the reliability of the LLMs-as-judges approach as leading models continue to evolve. Third, the influence of experiment parameters or alternate prompts were not investigated. In addition, our results based on a dataset rich in math and physics papers published in the past may not generalize well to papers in other scientific domains or future papers. We plan to address this limitation in future work. Last, since we were unable to recruit domain experts for manual annotation, it remains possible that some cases in the dataset contain problems that could not be detected based on the manuscript alone, or that the authors gave inaccurate reasons for retraction, which may pose an upper limit for model performance.
We noticed a concurrent work uploaded to arXiv on May 17 by Son et al. (2025), which also utilized the WithdrarXiv dataset for LLM-driven automatic error detection in scientific papers. Unlike their work which focused mainly on benchmarking LLMs, our work approached the same topic more from an application perspective. Their main metric overlaps substantially with our hit rate, although there is a difference in the meaning of $k$. Other key differences include: (1) Son et al. present a small but better annotated dataset designed solely for benchmarking, whereas our dataset is much larger, including a training set which may directly benefit future work; (2) they normalized the papers using OCR, whereas we assessed LLMs with PDF and LaTeX; (3) they used GPT-4.1 to determine whether an LLM’s error submissions match the gold error annotation, whereas we adopted multiple reasoning LLMs as judges. The strengths of their study include the use of domain experts for further quality control of their benchmark, and the use of resampling to study variations in LLM outputs. The two efforts are complementary, and both report the leading performance of o3 on scientific verification.
Data Availability
We will publish the WithdrarXiv-Check dataset, model outputs (including thinking outputs if available), and core experiment code to support future work by the research community. We will update this paper with the link to our public repository in the near future.
Acknowledgments
We thank community members of the Black Spatula Project for their inspiring discussions in December 2024, especially Ugo Dos Reis for highlighting the WithdrarXiv dataset, and Delip Rao for instructions on how to use this dataset. The authors also thank Yuxin Xu for her help with the API.
We welcome discussions, collaborations, and funding to further improve this work.
Appendix A Prompt Templates
Prompt for LLM quality checkers:
Please check the attached paper for critical errors and unsoundness problems that would invalidate the conclusions. You can ignore minor issues (e.g, typos and formatting errors) and limitations that have been properly acknowledged.
In your final output, give me up to {k} most critical problems as a JSON object using the following schema: Entry = {"Problem": str, "Location": str, "Explanation": str}, Return: list[Entry]. For location, give page number, section number, equation number, or whatever applicable. You can end the list early if there are fewer problems. No need to provide references.
{paper_attachment_or_LaTeX_script}
Prompt for LLM judges to determine hits:
My colleague was reading a paper and said there is a problem in it, as described below:
Problem: {problem}
Location: {location}
Explanation: {explanation}
I checked the paper and noticed that the authors have the following retraction comment:
{retraction_comment}
Did my colleague find exactly the same problem that the authors mentioned in the retraction comment? In your final answer, simply give "Yes" or "No". Make your best decision if you are unsure. You may explain your decision but please be concise.
Checker Model Judge Model
Gemini 2.5 Pro o3 (med)
PDF as an attachment
Gemini 2.5 Pro 64.9 62.4
Gemini 2.5 Flash 62.0 64.1
o3 (medium) 72.7 75.5
o4-mini (medium) 69.0 68.2
Claude 3.7 Sonnet 21.6 19.2
LaTeX script in prompt
Gemini 2.5 Pro 58.8 65.3
Gemini 2.5 Flash 58.8 61.6
o3 (medium) 75.5 80.4
o4-mini (medium) 66.1 71.4
Claude 3.7 Sonnet 44.5 40.8
Table 3: Hit rate results by individual judge.
Appendix B Model Parameters
Gemini 2.5 Pro and Gemini 2.5 Flash:
thinking budget: default (automatic)
tools: []
temperature: 0
seed: 42
o3 and o4-mini:
reasoning effort: defaults to "medium"
reasoning summary: "detailed"
tools: []
temperature and seed: not supported
Claude 3.7 Sonnet:
max tokens: 16,000
thinking type: "enabled"
thinking budget: 14,000
tools: []
temperature: 1 (required for thinking)
seed: not supported